Suppose that you are dealt a hand of 4 cards from a standard deck of 52 playing cards.
a.) What is the probability that you are dealt no more than 2 hearts?
b.) What is the probability that you are dealt no spades and exactly 1 eight?

Answer by stanbon(75887)

Suppose that you are dealt a hand of 4 cards from a standard deck of 52 playing cards.
a.) What is the probability that you are dealt no more than 2 hearts?
#( 0 hearts) = 13C0*39C4 = 39C4
#( 1 heart) = 13C1*39C3 = 13*39C3
#( 2 hearts) = 13C2*39C2
Ans:: (39C4 + 13*39C3 + 13C2*39C2+)/52C4
==============================================
b.) What is the probability that you are dealt no spades and exactly 1 eight?
#(no spades and 1 eight) = 3C1*36C3
P(no spades and 1 eight) = [3*36C3]/53C4
---------------------
Cheers,
Stan H.
-------------